16 November, 2016


The call for prayer of Isha marks the end of the twilight and the start of night. Just opened the roof of the observatory.

Tonight's targets are: V0725 Tau Photometric Data and some dark frames at -30C. CCD is at -25C right now.. still cooling down.

Water reservoir is at 19C ; i have set the cooling point to 13C.


50 Dark Frames of 120 seconds each have been acquired. 

V0725 Tau is still at 27 degrees altitude (very low for any scientific data acquisition) in the eastern horizon.


Focused the telescope with photometric I filter. Autoguiding at a 10th magnitude star at 0.5 Hz with the combination of Adaptive Optics and Mount autoguiding capability.

90 Seconds of exposure still saturated the star.


Changed it to 30 seconds now.. nope, still too much light! Lowered it to 15 seconds now and voila! a very nice 40K counts.

Acquisition begins now!


Have got 162 images in total. Now taking 15 sec Darks for some time now.


Reduced the data and performed photometry analysis in Maxim DL. As it turned out V0725 Tau star has right now 7.855 magnitude in I band. Submitted the data to AAVSO.